# Effective Game-Theoretic Motion Planning via Nested Search

### v1

@@ Abstract
To facilitate effective, safe deployment in the real world, individual robots must reason about interactions with other agents, which often occur without explicit communication. Recent work has identified game theory, particularly the concept of Nash Equilibrium (NE), as a key enabler for behavior-aware decision-making. Yet, existing work falls short of fully unleashing the power of game-theoretic reasoning. Specifically, popular optimization-based methods require simplified robot dynamics and tend to get trapped in local minima due to convexification. Other works that rely on payoff matrices suffer from poor scalability due to the explicit enumeration of all possible trajectories. To bridge this gap, we introduce Game-Theoretic Nested Search (GTNS), a novel, scalable, and provably-correct approach for computing NEs in general dynamical systems. GTNS efficiently searches the action space of all agents involved, while discarding trajectories that violate the NE constraint (no unilateral deviation) through an inner search over a lower-dimensional space. Our algorithm enables explicit selection among equilibria by utilizing a user-specified global objective, thereby capturing a rich set of realistic interactions. We demonstrate the approach on a variety of autonomous driving and racing scenarios where we achieve solutions in mere seconds on commodity hardware.
@@ 1 Introduction
Modern robotic applications—from autonomous driving to assistive robotics—require behavior-aware decision-making that accounts for interactions between the robot (the ego agent) and surrounding agents. These interactions often occur without explicit communication. Yet, to accomplish its mission effectively, the robot must reason about the behavior of other agents and the ways they respond to its actions. To address this challenge, this work seeks to advance the integration of game-theoretic reasoning into multi-robot decision-making by developing mechanisms that are both theoretically grounded and computationally efficient.
@@
The majority of approaches for behavior-aware decision-making [rudenko2020human] consider a two-step process where the ego robot first predicts the actions of surrounding agents [lee2017desire, salzmann2020trajectron++, jiang2023motiondiffuser] and then computes a best-response trajectory that considers the agents as having fixed routes, or by reasoning over the probabilities of taking various routes [schmerling2018multimodal]. Such approaches overlook the fact that agents can update their actions based on the behavior of other agents, which can result in overly conservative solutions. In some cases, it can give rise to the “frozen robot” problem, wherein the robot deems all possible actions as overly unsafe and thus chooses to take no action.
@@
More recent approaches aim to model interactions between agents through a game-theoretic perspective, thereby assuming that agents are rational and optimize a private objective function (e.g., travel time, safety, lead over opponent), while accounting for the fact that other agents are involved in the same process. For instance, some recent papers model interactions via a Stackelberg equilibrium and develop approaches based on dynamic programming to approximate the equilibrium [fisac2019hierarchical]. However, Stackelberg games give an unfair advantage to one of the players [le2022algames], and hence lead to less realistic behavior. In the context of autonomous driving, such approaches can create overly aggressive maneuvers for one of the vehicles and be overly conservative for the rest.
More recent work is focused on modeling interactions via a Nash Equilibrium (NE) [nash1950equilibrium], which does not create an artificial advantage between agents. A NE implies that no agent is better off by unilaterally deviating from the current solution. Although the game theory community has extensively studied NE for many years [nisan2007agt, shoham2009mas], most computationally-efficient results consider overly simplistic models that fail to capture the intricacies of real robotic systems, including continuous action and state spaces with non-convex constraints, as well as dependencies between the perception, planning, and control autonomy modules. For instance, some papers have attempted to approximate an NE using relaxations and utilizing differential dynamic programming [fridovich2020efficient], albeit with limited applicability to robotics due to the inability to capture collision-avoidance constraints.
@@
More specialized approaches that incorporate hard collision constraints and robot dynamics have been recently introduced in the robotics community, where NE is approximated through optimization-based approaches [wang2021game, le2022algames, spica2020real, le2021lucidgames], albeit with limited theoretical justification. For example, [le2022algames] utilizes a first-order approximation and, as such, cannot guarantee finding a global NE except in some special convex settings, which rarely occur in robotics. Specifically, the presence of obstacles, nonlinear dynamics, and multiple agents gives rise to multiple homotopy classes of NE solutions, which cannot be captured in a convex approximation. The authors of [wang2021game] suggest an iterative approach, which relies on [spica2020real], but that falls short of the sufficient conditions for NE convergence, attaining only the necessary conditions. This is due to a lack of convexity in both cases, caused by “holes” (dynamic obstacles) in the feasible convex configuration space, induced by inter-robot collision constraints.
@@
A different work [liniger2019noncooperative] considers a mathematically principled approach for efficiently obtaining a NE from a set of precomputed robot paths, under special conditions, such as right-of-way, which benefits one of the agents. Unfortunately, in a general setting, the approach boils down to a brute-force computation of the NE via the full payoff matrix, which is exponential in the number of robots. Moreover, the approach requires the explicit enumeration of all robot trajectories used, which can be costly if a rich set of strategies needs to be considered. A recent work [tikna2023graphs] invokes the bi-matrix game reasoning of [liniger2019noncooperative] in a receding horizon fashion to choose the immediate next action in a Stackleberg game, where the robots are confined to motions on a precomputed quasi-kinodynamic graph where edges represent Euler curves. Unfortunately, they provide no Equilibrium guarantees, and the use of Euler curves is highly problem- and model-specific.
@@
Learning-based approaches have gained significant popularity in robotics in recent years and have been successfully applied in behavior-aware settings in the context of head-to-head racing in computer games [wurman2022outracing], regulating mixed traffic flow [wu2021flow], and autonomous driving [wang2025generativeaiautonomousdriving], to name a few examples. However, as is typically the case with learning methods, they offer no explainability or generalizability and heavily rely on past experience for making correct decisions. Thus, their reliability remains unclear, especially in long tail events encompassing complex interactions between multiple agents [wang2025generativeaiautonomousdriving].
To summarize, existing work in game-theoretic motion planning lacks sufficient theoretical grounding or practical scalability.
@@
Contribution. We introduce an efficient and provably correct method Game-Theoretic Nested Search (GTNS) for computing a global NE for game-theoretic motion planning. Unlike aforementioned approaches in robotics, our method guarantees convergence to an NE without restricting the problem domain (e.g., convexity) and applies to general objective functions and dynamical constraints of individual robots. Furthermore, the approach enables the explicit tuning of the type of NE solution retrieved via a user-specified global objective function, ensuring that the returned NE minimizes this function among all available NEs. Examples of such global objectives include the sum of personal costs across all agents, or the cost of a specific competing agent.
@@
At the core of our methodology is a nested-search approach consisting of an outer and inner search method. The outer search implicitly explores a tree of potential multi-robot trajectories, while the inner search is invoked to verify that a given solution satisfies the NE requirement by considering improving unilateral deviations of individual robots. This search-based approach allows for an implicit encoding of trajectory-based constraints and incremental reuse of information, which is otherwise difficult to realize efficiently within an optimization-based approach [9922600] or in methods that require explicit enumeration of all possible trajectories [liniger2019noncooperative, tikna2023graphs].
@@
For computational efficiency, our nested-search approach implicitly explores the joint state space of the multi-robot system via graphs that efficiently encode the trajectory space of individual robots while adhering to the robots’ full dynamical constraints. We demonstrate the approach on a varierty of realistic autonomous driving and racing scenarios (including intersections, lane merges, following distance, opposing-lane passes, and racetrack overtakes) where we achieve solutions in mere seconds on commodity hardware. From the theory side, our algorithmic framework leads to a rigorous but natural proof for the attainment of an NE.
@@
(a) $\alpha=0.00,\ \lambda_{\mathrm{prox}}=0.000$
(b) $\alpha=0.25,\ \lambda_{\mathrm{prox}}=0.625$
(c) $\alpha=0.75,\ \lambda_{\mathrm{prox}}=1.875$
(d) $\alpha=1.00,\ \lambda_{\mathrm{prox}}=2.500$
[…]
Organization. This paper is organized as follows: Section 2 details the preliminaries of kinodynamic multi-robot motion planning and the game-theoretic concepts we employ, as well as declaring our problem statement and assumptions. In prepartion of the search algorithm, Section 3 considers a graph-based forumulation of the aforementioned problem. Section 4 describes in detail our algorithm GTNS, whose theoretical guarantees are then discussed in Section 5. Finally, Section 6 details our experimental setup and results, highlighting the novel aspects of our algorithm and its running time.
@@ 2 Preliminaries
We begin by introducing the notation for kinodynamic multi-robot motion planning, then formalize the underlying game-theoretic planning problem.
Consider $m\geq 2$ robots indexed by $i\in[m]:=\{1,\dots,m\}$. Each robot $i$ has a state space $\mathcal{X}^{i}\subseteq\mathbb{R}^{d_{i}}$ and a control space $\mathcal{U}^{i}\subseteq\mathbb{R}^{D_{i}}$. Its motion is governed by the dynamics
@@
Next, we consider robot trajectories resulting from integrating Eq. (1) for a given control law. For a given planning horizon $T>0$, let $\mathcal{U}^{i}_{T}$ be the set of admissible control functions $u^{i}:[0,T]\mapsto\mathcal{U}^{i}$ for robot $i$. Given an initial state $x_{0}^{i}\in\mathcal{X}^{i}$ and a control function $u^{i}\in\mathcal{U}^{i}_{T}$, the induced trajectory is the mapping $\pi^{i}_{x_{0}^{i},u^{i}}:[0,T]\mapsto\mathcal{X}^{i},$ where $\pi^{i}_{x_{0}^{i},u^{i}}(\cdot)$ is a solution to Equation (1) with $\pi^{i}_{x_{0}^{i},u^{i}}(0)=x_{0}^{i}$. The set of feasible trajectories for robot $i$ is
$\Pi^{i}_{T}:=\{\,\pi^{i}_{x_{0}^{i},u^{i}}\mid x_{0}^{i}\in\mathcal{X}^{i},\;u^{i}\in\mathcal{U}^{i}_{T}\,\}.$
@@
Each robot $i$ evaluates its performance via a cost functional
$J^{i}:\Pi^{i}_{T}\times\Pi^{-i}_{T}\mapsto\mathbb{R}_{\geq 0},$
where $\Pi^{-i}:=\times_{j\neq i}\Pi^{j}$ is the set of trajectories of the remaining robots with indices $m]\setminus\{i\}$. For trajectories $\pi^{i}\in\Pi^{i}_{T}$ and $\pi^{-i}\in\Pi^{-i}_{T}$,
$J^{i}(\pi^{i},\pi^{-i}):=\int_{0}^{T}c^{i}\bigl(\pi^{i}(t),\pi^{-i}(t)\bigr)\,\mathrm{d}t,$ (2)
with stage cost $c^{i}$ depending on states and controls (e.g., control effort, distance to obstacles, proximity to others).
@@ 2.1 Problem Definition
In this work, we seek to compute joint trajectories that satisfy the following requirement.
@@
A joint trajectory $\pi=(\pi^{1},\dots,\pi^{m})\in\Pi_{T}$ is a (pure) Nash equilibrium (NE) if, for every robot $i\in[m]$ and for all alternative feasible trajectories $\tilde{\pi}^{i}\in\Pi^{i}_{T}$ (from the same start state), no robot can reduce its cost by deviating while the other robots fix their trajectories, i.e.,
$J^{i}(\pi^{i},\pi^{-i})\;\leq\;J^{i}(\tilde{\pi}^{i},\pi^{-i}).$ (3)
Moreover, we are interested not only in finding any NE, but in selecting one that minimizes a user-defined global cost.
@@
$J(\pi):=\int_{0}^{T}c\bigl(\pi^{1}(t),\ldots,\pi^{m}(t)\bigr)\,\mathrm{d}t,\quad\pi\in\Pi_{T},$ (4)
with stage cost $c$ depending on joint states and controls.
We wish to find a joint trajectory $\pi^{*}\in\Pi_{T}$ (and its corresponding control laws) such that:
  1. 1.
    $\pi^{*}$ is a Nash equilibrium (Def. 1);
  2. 2.
    $\pi^{*}$ minimizes the global cost $J(\pi^{*})$ among all NEs;
  3. 3.
    the final state of $\pi^{*}$ is in the joint goal region, i.e., $\pi^{*}(T)\in\mathcal{X}_{\mathrm{goal}}:=\times_{i=1}^{m}\mathcal{X}^{i}_{\mathrm{goal}}$;
  4. 4.
    $\pi(t)\in\mathcal{X}_{f}$ for all $t\in[0,T]$.
@@
Discussion. A few comments are in order. We consider an open-loop setting: an NE solution is computed once for the horizon $T$ and then executed, rather than replanned at every time step. This setting is common in robotics (see, e.g., [ZhuB23]), motivated by the high computational demands of computing NEs, which currently preclude real-time feedback control. The resulting NE solutions promote stable execution, since local controllers cannot improve their cost by deviating from the prescribed trajectory.
Our formulation naturally extends to the $\varepsilon$–NE setting [TsaknakisS08], in which no robot can improve its cost by more than a factor of $(1+\varepsilon)$ through unilateral deviation, where Eq. (3) is replaced with the inequality
$J^{i}(\pi^{i},\pi^{-i})\;\leq\;(1+\varepsilon)\,J^{i}(\tilde{\pi}^{i},\pi^{-i}).$ (5)
@@
We emphasize that we seek global NEs, where robots may choose any action in their control spaces This contrasts with local NEs, which are implicitly considered in, e.g., [le2022algames, wang2021game, spica2020real, rowold2024open, fridovich2020efficient], where the solution is restricted to a subset of the joint state space (e.g. due to linearization and convexification). Local NEs may not correspond to any global NE, leading to unstable or otherwise poorly chosen execution. See illustration in Appendix 0.A, Fig. 7.
Another key difference from prior work is our requirement that the chosen NE optimize a global objective. This promotes predictability and explainability, as (under mild conditions) a unique global optimum exists, ensuring all robots converge to the same NE even if using different algorithms. Without this constraint, agents may independently converge to distinct NEs that are not jointly valid, a source of instability noted in [wang2021game, le2022algames], and tested in Section 6.
This framework can accommodate rich multi-agent preferences (e.g., effort $\int\!\|u\|^{2}$, comfort/jerk, rule compliance, barrier margins, connectivity/formation, visibility) via the user-defined objective, and in fact, even enables targeting specific equilibria: Let $\pi$ be a target NE that yields per-agent costs $g^{i}:=J^{i}(\pi^{i})$, by choosing $J=\sum_{i}\bigl|J^{i}-g^{i}\bigr|$ as our global cost function, our algorithm finds the targeted NE $\pi$ (upto tie-breaking between NEs with the exact same $g^{i}$s).
@@
A remaining challenge is incentivizing robots to follow the chosen solution. An NE optimizing $J$ may favor certain robots, motivating others to defect. In cooperative domains (e.g., driving), $J$ can encode social welfare or fairness and may even be set by an external arbitrator (e.g., to resolve right-of-way). In competitive settings, the more general solution concept of correlated equilibrium (CE) [aumann1974subjectivity] could be applied: a mediator samples from a distribution and privately recommends actions. Every NE is a CE, but CEs enable coordination on outcomes that are more efficient or fair. In our setting, multiple solutions to Problem 1 could be combined into a CE by sampling between them with carefully chosen probabilities. This motivates extending our planner: instead of selecting a single NE, output a distribution over equilibria that all robots are incentivized to follow, since CEs guarantee no profitable unilateral deviations in expectation.
@@
Figure 2: Kinodynamic grid graph. Each discrete state $(x,y,\theta,v,\delta)$ is drawn with two arrows anchored at $(x,y)$: a blue heading arrow pointing in direction $\theta$ (fixed short length), and a red steering arrow pointing in direction $\theta+\delta$ with length $\propto v$. Graph edges are gray; a randomly selected subset is emphasized and color-coded by the source state.

### v2

@@ Abstract
To facilitate effective and safe deployment, individual robots must reason about interactions with other agents, which often occur without explicit communication. Recent work has identified game theory, particularly the concept of Nash Equilibrium (NE), as a key enabler for behavior-aware motion planning. Yet, existing work falls short of fully unleashing the power of game-theoretic reasoning. Specifically, optimization-based methods require simplified robot dynamics and may get trapped in local minima due to convexification. Other works that rely on the explicit enumeration of all possible trajectories suffer from poor scalability. To bridge this gap, we introduce Game-Theoretic Nested Search (GTNS), a scalable, and provably-correct approach for computing NEs in general dynamical systems. GTNS efficiently searches the action space of all agents involved, while discarding trajectories that violate the NE constraint through an inner search over a lower-dimensional space. Our algorithm enables explicit selection among equilibria by utilizing a user-specified global objective, thereby capturing a rich set of realistic interactions. We demonstrate the approach across a variety of autonomous driving and racing scenarios, achieving solutions in mere seconds on commodity hardware.
@@ 1 Introduction
Modern robotic applications—from autonomous driving to assistive robotics—require behavior-aware motion planning that accounts for interactions between the ego robot and surrounding agents. These interactions often occur without explicit communication. Yet, to accomplish its mission effectively, the robot must reason about the behavior of other agents and the ways they respond to its actions. To address this challenge, this work seeks to develop game-theoretic motion planners that are both theoretically grounded and computationally efficient.
@@
The majority of approaches for behavior-aware motion planning consider a two-step process where the ego robot computes a best-response trajectory after predicting the actions of surrounding agents [rudenko2020human, lee2017desire, salzmann2020trajectron++, jiang2023motiondiffuser, schmerling2018multimodal]. Such approaches overlook that agents can update their actions based on other agents’ behavior, leading to overly conservative solutions.
@@
More recent methods aim to model interactions between agents through a game-theoretic perspective, thereby assuming that agents are rational and optimize a private objective function (e.g., travel time, safety, lead over opponent), while accounting for the fact that other agents are involved in the same process. For instance, some papers model interactions via a Stackelberg equilibrium and develop approaches based on dynamic programming to approximate the equilibrium [fisac2019hierarchical]. However, Stackelberg games give an unfair advantage to one of the players [le2022algames], and hence lead to less realistic behavior. In the context of autonomous driving, such approaches can create overly aggressive maneuvers for one of the vehicles and be overly conservative for the rest.
@@
More advanced methods model interactions via a Nash equilibrium (NE) [nash1950equilibrium], wherein no agent is better off by unilaterally deviating from the current solution. Although the game-theory community has extensively studied NE [nisan2007agt, shoham2009mas], most computationally-efficient results consider overly simplistic models that fail to capture the intricacies of real robotic systems, including continuous action and state spaces with non-convex constraints, and long-horizon strategies. Some papers have attempted to approximate an NE using relaxations and utilizing differential dynamic programming [fridovich2020efficient], albeit with limited applicability to robotics due to the inability to capture collision-avoidance constraints. More specialized approaches that incorporate hard collision constraints and robot dynamics approximate NEs through optimization-based approaches [wang2021game, le2022algames, spica2020real, le2021lucidgames], albeit with limited theoretical justification. For example, [le2022algames] utilizes a first-order approximation and, as such, cannot guarantee finding a global NE except in some special convex settings, which rarely occur in robotics. Specifically, the presence of obstacles, nonlinear dynamics, and multiple agents gives rise to multiple homotopy classes of NE solutions, which cannot be captured in a convex approximation. An iterative approach [wang2021game], which relies on [spica2020real], falls short of the sufficient conditions for NE convergence, attaining only the necessary conditions. This is due to a lack of convexity in both cases, which follows from nonlinear dynamics and a non-convex state space.
@@
A different work [liniger2019noncooperative] considers a mathematically principled approach for efficiently obtaining a NE from a set of precomputed robot paths, under special conditions, such as right-of-way, which benefits one of the agents. Unfortunately, in a general setting, the approach boils down to a brute-force computation of the NE via the full payoff matrix, which is exponential in the number of robots. Moreover, the approach requires the explicit enumeration of all robot trajectories used, which can be costly if a rich set of strategies needs to be considered. A recent work [tikna2023graphs] invokes the bi-matrix game reasoning of [liniger2019noncooperative] in a receding horizon fashion to choose the immediate next action in a Stackelberg game, where the robots are confined to motions on a precomputed quasi-kinodynamic graph where edges represent Euler curves. Unfortunately, no equilibrium guarantees are given, and the use of Euler curves is highly problem- and model-specific.
An approach by LaValle and Hutchinson [lavalle2002optimal] also leverages graph search, explicitly constructing a frontier of all non-Pareto-dominated NE candidate trajectories for each node. While this yields equilibrium guarantees, it incurs a substantial computational burden due to the need to verify and maintain all such equilibria throughout the search.
@@
Learning-based approaches have been applied in behavior-aware settings in the context of head-to-head racing in computer games [wurman2022outracing], regulating mixed traffic flow [wu2021flow], and autonomous driving [wang2025generativeaiautonomousdriving], to name a few examples. However, as is typically the case with learning methods, they offer no explainability or generalizability and heavily rely on past experience for making correct decisions. Thus, their reliability remains unclear, especially in long tail events encompassing complex interactions between multiple agents [wang2025generativeaiautonomousdriving].
[…]
Figure 1: Multi-lane merge: Robot 1 (blue) merges across Robot 2’s (orange) lane into Robot 3’s (green) lane. Existing algorithms typically restricted to a single homotopy class, whereas our method—via tuning of the global objective (here, decreasing Robot 1’s priority (larger $\alpha$ in $J{=}(1{-}\alpha)J^{1}{+}\alpha(J^{2}{+}J^{3})$) and increasing proximity penalty weight $\lambda_{\mathrm{prox}}$)—yields distinct NE behaviors: (a) aggressive zip-merge; (b) zip-merge; (c) yield-then-merge; (d) over-cautious.
@@
Contribution. We introduce an efficient and provably correct method, termed Game-Theoretic Nested Search (GTNS), for computing NEs in motion planning. Our method guarantees convergence to an NE without restricting the problem domain (e.g., convexity) and applies to general objective functions and dynamical constraints of individual robots. Furthermore, the approach enables the explicit tuning of the NE solution retrieved via a user-specified global objective function, ensuring that the returned NE minimizes this function among all available NEs (see, e.g., Fig. 1).
@@
At the core of our methodology is a nested-search approach consisting of an outer and inner search method. The outer search implicitly explores a tree of potential multi-robot trajectories, while the inner search is invoked to verify that a given solution satisfies the NE requirement by considering improving unilateral deviations of individual robots. For computational efficiency, our nested-search approach implicitly explores the joint state space of the multi-robot system via graphs that efficiently encode the trajectory space of individual robots while adhering to the robots’ full dynamical constraints.
This search-based approach allows for an implicit encoding of trajectory-based constraints and incremental reuse of information, which is otherwise difficult to realize efficiently within an optimization-based approach [9922600] or in methods that require explicit enumeration of all possible trajectories [liniger2019noncooperative, tikna2023graphs].Our algorithm has conceptual similarities to the LaValle-Hutchinson approach [lavalle2002optimal], but crucially does not explicitly maintain all candidate NE trajectories, and employs a powerful pruning mechanism to discard non-NE trajectories (and their subtrees), rendering the search over complex and high-dimensional spaces tractable.
@@
We demonstrate the approach on a variety of realistic autonomous driving and racing scenarios (including intersections, lane merges, following distance, opposing-lane passes, and racetrack overtakes) where we achieve solutions in mere seconds on commodity hardware. From the theory side, our algorithmic framework leads to a rigorous but natural proof for the attainment of an NE.
@@
Organization. Sec. 2 provides preliminaries and our problem statement. Sec. 3 considers a graph-based formulation of the above problem. Sec. 4 describes our GTNS algorithm, whose theoretical guarantees are discussed in Sec. 5. Sec. 6 details experimental results. Conclusion and future work appears in Sec. 7.
@@ 2 Preliminaries
We begin by introducing the notation for kinodynamic multi-robot motion planning, then formalize the underlying game-theoretic planning problem. Consider $m\geq 2$ robots indexed by $i\in[m]:=\{1,\dots,m\}$. Each robot $i$ has a state space $\mathcal{X}^{i}\subseteq\mathbb{R}^{d_{i}}$ and a control space $\mathcal{U}^{i}\subseteq\mathbb{R}^{D_{i}}$. Its motion is governed by the dynamics
@@
For a given planning horizon $T>0$, let $\mathcal{U}^{i}_{T}$ be the set of admissible control functions $u^{i}:[0,T]\mapsto\mathcal{U}^{i}$ for robot $i$. Given an initial state $x_{0}^{i}\in\mathcal{X}^{i}$ and a control function $u^{i}\in\mathcal{U}^{i}_{T}$, the induced trajectory is the mapping $\pi^{i}_{x_{0}^{i},u^{i}}:[0,T]\mapsto\mathcal{X}^{i},$ where $\pi^{i}_{x_{0}^{i},u^{i}}(\cdot)$ is a solution to Eq. (1) with $\pi^{i}_{x_{0}^{i},u^{i}}(0)=x_{0}^{i}$. The set of feasible trajectories for robot $i$ is $\Pi^{i}_{T}:=\{\,\pi^{i}_{x_{0}^{i},u^{i}}\mid x_{0}^{i}\in\mathcal{X}^{i},\;u^{i}\in\mathcal{U}^{i}_{T}\,\}.$
@@
Each robot $i$ evaluates its performance via a cost functional $J^{i}:\Pi^{i}_{T}\times\Pi^{-i}_{T}\mapsto\mathbb{R}_{\geq 0},$ where $\Pi^{-i}:=\times_{j\neq i}\Pi^{j}$ is the set of trajectories of the remaining robots with indices $m]\setminus\{i\}$. For trajectories $\pi^{i}\in\Pi^{i}_{T}$ and $\pi^{-i}\in\Pi^{-i}_{T}$, we define
$\small J^{i}(\pi^{i},\pi^{-i}):=\int_{0}^{T}c^{i}\bigl(\pi^{i}(t),\pi^{-i}(t)\bigr)\,\mathrm{d}t,$} (2)
with stage cost $c^{i}$ depending on states and controls.
@@ 2.1 Problem Definition
We seek to compute joint trajectories that satisfy the following requirement.
@@
A joint trajectory $\pi=(\pi^{1},\dots,\pi^{m})\in\Pi_{T}$ is a (pure) Nash equilibrium (NE) if, for every robot $i\in[m]$ and for all alternative feasible trajectories $\tilde{\pi}^{i}\in\Pi^{i}_{T}$ (from the same start state), no robot can reduce its cost by deviating while the other robots fix their trajectories, i.e., $J^{i}(\pi^{i},\pi^{-i})\;\leq\;J^{i}(\tilde{\pi}^{i},\pi^{-i})$.
Moreover, we wish to compute an NE that minimizes a given global cost.
@@
$\small J(\pi):=\int_{0}^{T}c\bigl(\pi^{1}(t),\ldots,\pi^{m}(t)\bigr)\,\mathrm{d}t,\quad\pi\in\Pi_{T},$} (3)
with stage cost $c$ depending on joint states and controls. We wish to find a joint trajectory $\pi^{*}\in\Pi_{T}$ (and its corresponding control laws) such that: (1) $\pi^{*}$ is a Nash equilibrium (Def. 1) that minimizes the global cost $J(\pi^{*})$ among all NEs; (2) the final state of $\pi^{*}$ is in the joint goal region, i.e., $\pi^{*}(T)\in\mathcal{X}_{\mathrm{goal}}:=\times_{i=1}^{m}\mathcal{X}^{i}_{\mathrm{goal}}$; (3) $\pi(t)\in\mathcal{X}_{f}$ for all $t\in[0,T]$.
@@
Discussion. We consider an open-loop setting: an NE solution is computed once for a horizon and then executed, rather than replanned at every time step. This setting is common in robotics (see, e.g., [ZhuB23]), motivated by the high computational demands of computing NEs, which currently preclude real-time feedback control. The resulting NE solutions promote stable execution, since local controllers cannot improve their cost by deviating from the prescribed trajectory.
@@
We seek global NEs, where robots may choose any action in their control spaces This contrasts with local NEs, which are implicitly considered in, e.g., [le2022algames, wang2021game, spica2020real, rowold2024open, fridovich2020efficient], where the solution is restricted to a subset of the joint state space (e.g. due to linearization and convexification). Local NEs may not correspond to any global NE, leading to unstable or otherwise poorly chosen execution. See illustration in App. 0.B, Fig. 6. Another key difference from prior work is that we require the chosen NE to optimize a global objective. This promotes predictability and explainability, as (under mild conditions) a unique global optimum exists, ensuring all robots converge to the same NE.
@@
A remaining challenge is incentivizing robots to follow the chosen solution. An NE optimizing a specific cost function $J$ may favor certain robots, motivating others to defect. In cooperative domains (e.g., driving), $J$ can encode social welfare or fairness and may even be set by an external arbitrator (e.g., to resolve right-of-way). In competitive settings, the more general solution concept of correlated equilibrium (CE) [aumann1974subjectivity] could be applied: a mediator samples from a distribution and privately recommends actions, which can enable coordination on outcomes that are more efficient or fair. We leave this for future work.
@@
Figure 2: Kinodynamic grid graph. At each $(x,y)$ lattice coordinate (blue dots) there are various $\theta,\ v$ and $\delta$ values. Graph edges are gray; a randomly selected subset is emphasized and colored by the source state.